# Weyl’s Quantifiers

### v2

@@ 1 Introduction
Some likely historical sources for Weyl’s interpretation are Schlick’s 1918 Allgemeine Erkenntnislehre and Alexander Pfänder’s 1919 Logik (Sundholm 1994, 125sq), or Schopenhauer’s 1859 Die Welt als Wille und Vortstellung (Detlefsen 2011, 90). It is entirely reasonable to believe that Weyl was fully aware of these works. It seems no less reasonable to assume that Weyl was also acquainted with Husserl’s phenomenology of values, and in particular his lectures on axiology (Husserl 1988), delivered in the winter semester 1908/09 and the summer semesters 1911 and 1914 in Göttingen, where Weyl was based until he took up his position at the ETH in Zurich, in 1913. Despite McCarty’s reservations, I think it would be very important and extremely interesting to investigate Husserlian axiology as a basis for Weyl’s interpretation of the logical signs. But this historical investigation will have to be pursued elsewhere. Here I want to focus on Weyl’s interpretation, independently of any sources that might have helped him to develop it. [^3] For discussion of Husserlian axiology, see Mulligan 2004. Weyl was certainly familiar with the Logische Untersuchungen, where Husserl gives a preliminary account of values (see Mulligan 2022), but it is an open question if that would have been sufficient as a basis for Weyl’s interpretation of the logical signs.
@@ 2 Against the Ramseyan view
The refutation of the Ramseyan view starts by clarifying a central element of Weyl’s philosophy of mathematics: the notion of proper judgment, or what he called judgment in a proper sense. As early as 1918, in his Das Kontinuum, he explained this notion as part of his characterization of what he called elementary inferences in mathematics: a proper judgment is a judgment that represents a state of affairs, and elementary inferences are inferences that involve only such judgments (Weyl 1918, 17). Weyl understood inferential correctness to depend essentially on representational correctness, and defined entailment in terms of an order relation between correctly represented states of affairs. For example, to borrow an illustrious example from Locke, reasoning from “AB is an arch of a circle” to “AB is less than the whole circle” is inferentially correct in virtue of the order relation, and more exactly the containment relation, between these two correctly represented states of affairs. According to Weyl, a genuine mathematical proof should deploy only inferentially correct reasoning and, thus, should contain only representationally correct judgments. He explicitly endorsed a correctness-first account of knowledge: “[representational] correctness … remains throughout the ultimate source of knowledge.” (op. cit., 11) What led him to this account was his early acquiescence to a phenomenological epistemology, which conceived of correctness as a norm, in a deontic sense: correct judgments are judgments that we ought to judge. For Weyl then, a genuine proof requires that we fully discharge our unconditional obligation to judge correctly. [^7] On this epistemological background, Weyl criticized Dedekind’s “perverted” conception of proof, which he (Weyl) thought promoted a notion of inferential correctness that is independent of representational correctness and, therefore, epistemologically problematic. For some discussion of this criticism, see Toader 2016. The obligation to judge correctly may be regarded as an epistemic obligation, although this notion typically refers to obligations in matters of belief or opinion, rather than judgment. For discussions of the notion of correctness, as understood in the phenomenological tradition, see Mulligan 2017, and Textor 2019, 2022. As I see it, Weyl’s notion of representational correctness comes down to, or at least includes, the requirement that mathematical reasoning should be conducted within a fully interpreted language.
@@
Weyl did not say that universal statements are meaningless, either. Just as above, one could consider them meaningless if one mistook them for proper judgments. But he again emphasized that they are not to be taken as proper judgments. What roles do universal statements play in mathematics then, on Weyl’s correctness-first account of mathematical knowledge? How can they serve the provision of genuine proof? Weyl wrote that universal statements are instructions for the recovery of proper judgments. As such, he considered them valuable insofar as they prompt us to act in certain ways – to recover proper judgments. Again, given Weyl’s emphasis our obligation to judge correctly, I think it is also natural to understand instructions normatively as generating a further obligation: we ought to recover correct judgments. This obligation is conditional on universal quantification. I will revisit this point as well in the next section, where I give a more exact reconstruction of Weyl’s reasoning in the second fragment, which will clearly show that the rules of Weyl’s universal quantifier are different than the usual logical rules.
@@
I turn now to explaining in more detail why I think that, according to Weyl, quantified statements generate obligations, and how this clarifies his interpretation of the logical signs and his actual reasons for rejecting the law of excluded middle. Of course, it’s difficult to say that the reconstruction I propose is entirely faithful to Weyl’s thought, but I think it comes closest, and in any case, it is preferable to the Ramseyan view.
@@ 3 Transfinite mathematics as a normative system
But there is a problem with claim 4’. Recovering a proper judgment whenever an object of the domain is given is not always possible, because we cannot run through an infinite set (Weyl 1921, 54). To address this point, Weyl proposed that a universal statement is to be regarded as a special kind of existential statement, one that asserts the existence of an essence. Thus, an instruction for recovering a correct judgment should be reconceived as an indication of the existence of an essence. This suggests the replacement of claims 4’ and 5’ by the following ones:
@@
This understanding of the function of quantified statements implies that Weyl’s quantifiers have different rules than the familiar rules of the usual logical quantifiers. These rules are expressed in claims 2’ and 4”. They may be regarded as deontic rules for quantifier elimination, rules that prescribe how we ought to act when we assert quantified statements. Reconstructed in this way, Weyl’s view is neither strange, nor really metaphorical. Unlike Hilbert, who understood quantified statements as partial assertions or ideal statements that may have a function in logic, Weyl assigned them the normative function just described. But if this reconstruction is accurate, as I believe it to be, then it follows that he never conceived of transfinite mathematics as a game with meaningless symbols, despite the fact that he did attribute this view (rightly or wrongly) to Hilbert. Rather, Weyl saw it as a normative extension of elementary mathematics, as a system of conditional obligations. [^8] Conditional obligations have been studied, of course, since at least the 1950s (see e.g. von Wright 1956), and various problems with both unconditional and conditional obligations, especially contrary-to-duty obligations, have been extensively discussed (see e.g. Chisholm 1963, van Fraassen 1972, and most recently Fine 2024). My reconstruction would no doubt be strengthened by solutions to at least some such problems as might arise for Weyl’s conception of transfinite mathematics, but this will have to be addressed in future work.
@@ 4 Nested quantifiers
Before we see how Weyl answers this question, we have to understand why he asks this question. Why must we be dealing here with an abstract from a judgment instruction, rather than with an instruction for recovering a judgment abstract? Key to this is his claim that universal generalizations of existential statements, like “For every number m there is a number n such that the relation R(m, n) holds between them.” must be interpreted as “There is a law $\phi$ such that for every number m the relation $R$ holds between $m$ and $\phi(m)$.” More generally, any instruction for recovering a judgment abstract must be interpreted as the indication of a law that backs the instruction for recovering a proper judgment. This reduces the case of universal generalizations of existential statements to the unproblematic case of existential generalizations of universal statements. Taking once more into account his correctness-first epistemology, my reconstruction explains what obligations are generated by universal generalizations of existential statements and why we value the latter:
@@
However, for precisely the epistemic reasons indicated by Weyl, it is doubtful that we can always discharge the obligation to recover an indication of the existence of a correct judgment. But then, without an indication of the existence of a correct judgment, the obligation to search for a correct judgment would not be generated, and therefore we would at least sometimes fail to discharge our obligation to judge correctly.
@@ 5 Conclusion
The Ramseyan view, which has been the predominant view of Weyl’s interpretation of the logical signs, conflates the latter with Hilbert’s. It maintains that they both had the same understanding of quantification, and they both demarcated meaningful from meaningless statements in one and the same way. In this paper, I have offered some reasons to resist this view. I have pointed out that a central element of Weyl’s interpretation is his correctness-first account of mathematical knowledge. This account, according to which we are under an unconditional obligation to judge correctly, suggests an understanding of quantified statements as generating further, conditional obligations to act in ways that expand the repository of correct judgments. I have argued that this justifies the claim that Weyl conceived of transfinite mathematics as a normative extension of elementary mathematics. This clarifies Weyl’s semantic reasons for rejecting the validity of the law of excluded middle, which have nothing to do with the reasons attributed to him by the Ramseyan view.
My reconstruction of Weyl’s interpretation is articulated by claims 1’, 2’, 3’, 4”, and 5”. I have also offered some preliminary thoughts on how to understand his remarks on nested quantification, and the obligations generated by statements with quantifiers within the range of other quantifiers, as articulated by claims 6, 7, 8, and 9. I think that the reconstruction can further be fruitfully used in an attempt to clarify Weyl’s remarks on choice sequences, and in particular, the fact, often noted, that for him, “the universal quantifier ranges over lawless, the existential over lawlike sequences” (van Atten, van Dalen, and Tieszen 2002, 220sq). If a universal statement cannot generate an obligation to recover a correct judgment, but only an obligation to search for a law in order to recover a correct judgment, then it follows, in particular, that universal quantification over lawless sequences cannot generate a obligation to recover a correct judgment about lawless sequences, but only an obligation to search for a law. What could be then recovered on the basis of a law is, of course, only a correct judgment about lawlike sequences. Note that this does not entail that there cannot be correct judgments about lawless sequences, but only that the universal quantifier does not generate any obligation to recover such judgments. More needs to be said, however, to provide a fully satisfactory account.
@@ Acknowledgments
Thanks to the editors for inviting me to contribute to this volume, to the referees for their comments, and to Kevin Mulligan for his thoughts on a previous lengthier draft of this paper. I am greatly indebted to the late David Charles McCarty for his support and learnedness, which he shared most kindly whenever we met.

### v3

@@ 1 Introduction
Some likely historical sources for Weyl’s interpretation are Schlick’s 1918 Allgemeine Erkenntnislehre and Alexander Pfänder’s 1919 Logik (Sundholm 1994, 125sq), or Schopenhauer’s 1859 Die Welt als Wille und Vortstellung (Detlefsen 2011, 90). It is entirely reasonable to believe that Weyl was fully aware of these works. It seems no less reasonable to assume that Weyl was also acquainted with Husserl’s phenomenology of values, and in particular his lectures on axiology (Husserl 1988), delivered in the winter semester 1908/09 and the summer semesters 1911 and 1914 in Göttingen, where Weyl was based until he took up his position at the ETH in Zurich, in 1913. Despite McCarty’s reservations, I think it would be very important and extremely interesting to investigate Husserlian axiology as a basis for Weyl’s interpretation of the logical signs. But this historical investigation will have to be pursued elsewhere. Here I want to focus on Weyl’s interpretation, quite independently of any sources that might have helped him to develop it. [^3] For discussion of Husserlian axiology, see Mulligan 2004. Weyl was certainly familiar with the Logische Untersuchungen, where Husserl gives a preliminary account of values (see Mulligan 2022), but it is an open question if that would have been sufficient as a basis for Weyl’s interpretation of the logical signs.
@@ 2 Against the Ramseyan view
The refutation of the Ramseyan view starts by clarifying a central element of Weyl’s philosophy of mathematics: the notion of proper judgment, or what he called judgment in a proper sense. As early as 1918, in his Das Kontinuum, he explained this notion as part of his characterization of what he called elementary inferences in mathematics: a proper judgment is a judgment that represents a state of affairs, and elementary inferences are inferences that involve only such judgments (Weyl 1918, 17). Weyl understood inferential correctness to depend essentially on representational correctness, and defined entailment in terms of an order relation between correctly represented states of affairs. For example, to borrow an illustrious example from Locke, reasoning from “AB is an arch of a circle” to “AB is less than the whole circle” is inferentially correct in virtue of the order relation, and more exactly the containment relation, between these two correctly represented states of affairs. According to Weyl, a genuine mathematical proof should deploy only inferentially correct reasoning and, thus, should contain only representationally correct judgments. He explicitly endorsed a correctness-first account of knowledge: “correctness … remains throughout the ultimate source of knowledge.” (op. cit., 11) As the term is used here, “correctness” designates an epistemic aspect of the representational capacity of judgments whose semantic aspect is truth. What led him to this account was his early acquiescence to a phenomenological epistemology, which conceived of correctness as a norm, in a deontic sense: correct judgments are judgments that we ought to judge. For Weyl then, a genuine proof requires that we fully discharge our unconditional obligation to judge correctly. [^7] On this epistemological background, Weyl criticized Dedekind’s conception of proof, which he (Weyl) thought promoted a notion of inferential correctness that is independent of representational correctness and, therefore, is epistemologically problematic. For some discussion of this criticism, see Toader 2016. The obligation to judge correctly may be regarded as an epistemic obligation, although this notion typically refers to obligations in matters of belief or opinion, rather than judgment. For discussions of the notion of correctness, as understood in the phenomenological tradition, see Mulligan 2017, and Textor 2019, 2022.
@@
Weyl did not say that universal statements are meaningless, either. Just as above, one could consider them meaningless if one mistook them for proper judgments. But he again emphasized that they are not to be taken as proper judgments. What roles do universal statements play in mathematics then, on Weyl’s correctness-first account of mathematical knowledge? How can they serve the provision of genuine proof? Weyl wrote that universal statements are instructions for the recovery of proper judgments. As such, he considered them valuable insofar as they prompt us to act in certain ways – to recover proper judgments. Again, given Weyl’s emphasis on our obligation to judge correctly, I think it is also natural to understand instructions normatively as generating a further obligation: we ought to recover correct judgments. This obligation is conditional on universal quantification. I will revisit this point as well in the next section, where I give a more exact reconstruction of Weyl’s reasoning in the second fragment, which will clearly show that the rules of Weyl’s universal quantifier are different than the usual logical rules.
@@
I turn now to explaining in more detail why I think that, according to Weyl, quantified statements generate obligations, and how this clarifies his interpretation of the logical signs and his actual reasons for rejecting the law of excluded middle. Of course, it’s difficult to say that the reconstruction I propose is entirely faithful to Weyl’s thought, but I think it comes close to it, and in any case, it is preferable to the Ramseyan view.
@@ 3 Transfinite mathematics as a normative system
But there is still a problem with claim 4’. Recovering a proper judgment whenever an object of the domain is given is not always possible, because we cannot run through an infinite set (Weyl 1921, 54). To address this point, Weyl proposed that a universal statement is to be regarded as a special kind of existential statement, one that asserts the existence of an essence. Thus, an instruction for recovering a correct judgment should be reconceived as an indication of the existence of an essence. This suggests the replacement of claims 4’ and 5’ by the following ones:
@@
This understanding of the function of quantified statements implies that Weyl’s quantifiers have different rules than the familiar rules of the usual logical quantifiers. These rules are expressed in claims 2’ and 4”. They may be regarded as deontic rules for quantifier elimination, rules that prescribe how we ought to act when we assert quantified statements. Reconstructed in this way, Weyl’s view is neither strange, nor really metaphorical. Unlike Hilbert, who understood quantified statements as partial assertions or ideal statements that may have a function in logic, Weyl assigned them the normative function just described. But if this reconstruction is accurate, as I believe it to be, then it follows that he never conceived of transfinite mathematics as a game with meaningless symbols, despite the fact that he did attribute this view (rightly or wrongly) to Hilbert. Rather, Weyl saw it as a normative extension of elementary mathematics, as a system of conditional obligations. [^8] Conditional obligations have been studied, of course, since at least the 1950s (see e.g. von Wright 1956), and various problems and puzzles with both unconditional and conditional obligations, especially contrary-to-duty obligations, have been extensively discussed (see e.g. Chisholm 1963, van Fraassen 1972, and most recently Fine 2024). My reconstruction would be strengthened by solutions to at least some such problems as might arise for Weyl’s conception of transfinite mathematics, but this will have to be addressed in future work.
@@ 4 Nested quantifiers
Before we see how Weyl answers this question, we have to understand why he asks this question. Why must we be dealing here with an abstract from a judgment instruction, rather than with an instruction for recovering a judgment abstract? Key to this is his claim that universal generalizations of existential statements, like “For every number m there is a number n such that the relation R(m, n) holds between them.” must be interpreted as “There is a law $\phi$ such that for every number m the relation $R$ holds between $m$ and $\phi(m)$.” More generally, any instruction for recovering a judgment abstract must be interpreted as the indication of a law that backs the instruction for recovering a proper judgment. This reduces the case of universal generalizations of existential statements to the unproblematic case of existential generalizations of universal statements. Taking once more into account Weyl’s correctness-first epistemology, my reconstruction explains what obligations are generated by universal generalizations of existential statements and why we value the latter:
@@
However, for precisely the epistemic reasons indicated by Weyl, it is doubtful that we can always discharge the obligation to recover an indication of the existence of a correct judgment. But then, without an indication of the existence of a correct judgment, the conditional obligation to search for a correct judgment would not be generated, and therefore we would at least sometimes fail to discharge our unconditional obligation to judge correctly.
@@ 5 Conclusion
The Ramseyan view, which has been the predominant view of Weyl’s interpretation of the logical signs, conflates the latter with Hilbert’s. It maintains that they both had the same understanding of quantification, and they both demarcated meaningful from meaningless statements in one and the same way. In this paper, I have offered some reasons to resist this view. I have pointed out that a central element of Weyl’s interpretation is his correctness-first account of mathematical knowledge. This account, according to which we are under an unconditional obligation to judge correctly, suggests an understanding of quantified statements as generating further, conditional obligations to act in ways that expand the repository of correct judgments. I have argued that this justifies the claim that Weyl conceived of transfinite mathematics as a normative extension of elementary mathematics. This clarifies Weyl’s semantic reasons for rejecting the validity of the law of excluded middle, which have nothing to do with the reasons attributed to him on the Ramseyan view.
My reconstruction of Weyl’s interpretation is articulated by claims 1’, 2’, 3’, 4”, and 5” above. I have also offered some preliminary thoughts on how to understand his remarks on nested quantification, and the obligations generated by statements with quantifiers within the range of other quantifiers, as articulated by claims 6, 7, 8, and 9. I think that the reconstruction can further be fruitfully used in an attempt to clarify Weyl’s remarks on choice sequences, and in particular, the fact, often noted, that for him, “the universal quantifier ranges over lawless, the existential over lawlike sequences” (van Atten, van Dalen, and Tieszen 2002, 220sq). If a universal statement cannot generate an obligation to recover a correct judgment, but only an obligation to search for a law in order to recover a correct judgment, then it follows, in particular, that universal quantification over lawless sequences cannot generate a obligation to recover a correct judgment about lawless sequences, but only an obligation to search for a law. What could be then recovered on the basis of a law is, of course, only a correct judgment about lawlike sequences. Note that this does not entail that there cannot be correct judgments about lawless sequences, but only that the universal quantifier does not generate any obligation to recover such judgments. More needs to be said, however, to provide a fully satisfactory account.
@@ Acknowledgments
Thanks to the editors for inviting me to contribute to this volume, to the referees for their comments, and to Kevin Mulligan for his thoughts on a previous lengthier draft of this paper. I am also indebted to the late David Charles McCarty for his encouragement and learnedness, which he always shared most kindly.